# A bottleneck model with shared autonomous vehicles: Scale economies and price regulations

### v1

@@ 4 First-best policy
So far, we have examined how the various pricing regulations influence commuters’ equilibrium costs. We now shift our attention to social costs that incorporate the service operator’s profit, and to transport policies designed to achieve an efficient transportation system.
This section examines the first-best policy that minimizes the social cost. Specifically, we assume that, in addition to price regulation, a policymaker can impose time-varying congestion tolls that completely internalize congestion externalities. The equilibrium condition is formulated under the first-best policy. We then derive the equilibrium flow and cost patterns under the first-best policy and analyze their properties.
@@ 4.1 Formulation
$\min_{\left\{n_{n}(t),n_{a}(t)\right\}_{t\in\mathbb{R}}}\quad$ $SC(\left\{n_{n}(t),n_{a}(t)\right\}_{t\in\mathbb{R}}):=\int_{t\in\mathbb{R}}\left[n_{n}(t)(t_{f}+s(t)+F_{n})+n_{a}(t)(\theta t_{f}+s(t)+m)\right]\mathrm{d}t+F_{a}$ (36)
s.t. $\int_{t\in\mathbb{R}}n_{n}(t)\mathrm{d}t+\int_{t\in\mathbb{R}}n_{a}(t)\mathrm{d}t=N,$ (37)
$n_{n}(t)+\kappa n_{a}(t)\leq\mu,\quad\quad\forall t\in\mathbb{R},$ (38)
$n_{n}(t)\geq 0,\quad n_{a}(t)\geq 0,\quad\quad\forall t\in\mathbb{R}.$ (39)
@@
We then consider a first-best policy that achieves this outcome in equilibrium. The policy consists of optimal price regulation together with a time-varying congestion toll. Let $\tau(t)$ denote the congestion toll paid by NV commuters who arrive at their destination at time $t$. We assume that each SAV commuter pays $\kappa\tau(t)$, because an SAV imposes congestion externalities that are $\kappa$ times those generated by an NV commuter. We further assume that price regulation follows MC pricing; that is, the fare equals the marginal cost mm. [^6] Charging differentiated congestion tolls by vehicle class is technically feasible and has already been implemented in practice. An alternative is to establish, outside the transport system, a market for tradable access rights or credits that allows commuters to pre-purchase bottleneck permits priced according to their vehicle classes. Under these assumptions, the commuting costs for NV and SAV commuters arriving at time $t$ are given by
@@
$c_{a}^{FB}(t)=s(t)+\theta t_{f}+m+\kappa\tau(t).$ (41)
Using them, the equilibrium conditions under the first-best policy are expressed as follows:
@@
It is straightforward to verify that the proposed policy, namely MC pricing combined with vehicle-specific congestion tolls, is first best by examining the optimality conditions of the optimization problem (36)–(39). The equilibrium conditions coincide with the first-order conditions when the Lagrange multiplier for constraint (37) is set to $c^{FB*}$ and the multiplier for constraint (38) is set to $\tau(t)$. Consequently, an equilibrium that satisfies the equilibrium conditions attains the first-best optimum, and the proposed policy is therefore first best.
@@ 4.2 Characterization of the equilibrium under the first-best policy
The results indicate that a flow pattern consisting exclusively of SAVs cannot arise under the first-best equilibrium. The reason lies in the condition $B>0$, which represents that the marginal cost of using an SAV exceeds that of an NV. Because queues are eliminated in the first-best equilibrium, commuters’ mode choice depends only on differences in tolls and marginal costs. During periods with low tolls (e.g., immediately after the bottleneck becomes active or just before it becomes inactive), the higher marginal cost of using an SAV outweighs the toll saving from choosing one. As a result, commuters have an incentive to choose NVs.
Figure 4: Time-varying congestion tolls in the first-best equilibrium. $N=10,000$, $\theta=0.5$, $\beta=0.3$, $\gamma=2.0$, $\mu=1.0$, $t_{f}=2.0$, $F_{n}=10,000$, $m=10500$.
Figure 4 illustrates the time-varying congestion tolls $\tau(t)$ for NV commuters and $\kappa\tau(t)$ for SAV commuters. In the figure, the tolls peak around the desired arrival time because SAVs are used during this period. An SAV commuter pays $\kappa\tau(t)$, which is $\kappa$ times the toll paid by an NV commuter and therefore lower. Taking this into account, the policymaker sets $\tau(t)$ sufficiently high to prevent queues from forming among SAV commuters. In addition, a comparison of $\tau(t)$ and $\kappa\tau(t)$ shows that their gap is large when SAVs are used and narrows as the destination-arrival time moves away from the desired arrival time. Consequently, in periods where this gap is small, commuters have an incentive to choose NVs, whose marginal cost is lower than that of SAVs.
@@ 4.3 Pareto-improvement condition
Using the closed-form solution, we examine whether the first-best policy can achieve a Pareto improvement. Relative to the equilibrium under any fare-setting scenario without congestion tolls, the first-best policy reduces social cost. It is desirable not only to reduce the social cost but also to improve the individual commuting cost. To identify a sufficient condition for this Pareto improvement, we compare the commuting cost $c^{FB*}$ in the first-best equilibrium with that under MC pricing $c^{MC*}$, which yields the lowest commuting cost among the three fare-setting scenarios.
For this analysis, the following capacity-VOT index $\eta$ plays a central role:
@@
whose value is determined by the relative sizes of $\kappa$ and $\theta$:
@@
This relationship shows that when the capacity effect dominates ($\eta>1$), $N_{a}^{FB*}$ exceeds $N_{a}^{MC*}$. The first-best policy, therefore, requires an additional increase in SAV ridership. As the number of SAV commuters increases, the total capacity effect grows accordingly. This results in reducing commuters’ schedule costs and thereby their commuting costs.
By contrast, when $\eta<1$, we see that the first-best policy works in the opposite direction and reduces the number of SAV commuters. As a result, the total capacity effect in the first-best equilibrium becomes smaller, and schedule costs can increase. The same phenomenon occurs in Case (ii) as well: since SAVs are not used in Case (ii), the first-best policy reduces the SAV ridership and thereby increases commuting costs. These findings imply that congestion tolls should be introduced taking into account the strength of the capacity effect, i.e., the technological maturity of SAVs.
@@ 4.4 Self-financing principle
We demonstrated that the first-best equilibrium can be achieved by combining MC pricing with a congestion toll. For a congestion-pricing scheme to be socially acceptable, it is desirable to redistribute the toll revenue to commuters. As a benchmark form of redistribution, we consider using the revenue to finance the expansion of bottleneck capacity. We then investigate whether the self-financing principle (Mohring and Harwitz, 1962) holds, namely, whether the revenue generated by the optimal congestion toll covers the cost of the optimal capacity level.
@@
$\begin{cases}\cfrac{dK(\mu)}{\partial\mu}-\int_{t\in\mathbb{R}}\tau(t)\mathrm{d}t=0\quad&\text{if}\quad\mu>0,\\ \cfrac{dK(\mu)}{\partial\mu}-\int_{t\in\mathbb{R}}\tau(t)\mathrm{d}t\geq 0\quad&\text{if}\quad\mu=0.\end{cases}$ (59)
@@ 5 Second-best policy
In this section, we analyze the second-best policy that minimizes the social cost when congestion externalities cannot be internalized. We consider a setting in which policymakers can regulate only the SAV fare. The fare is therefore set to minimize social cost in the equilibrium that satisfies both the departure-time and mode-choice conditions. We first derive the second-best fare and the corresponding number of SAV commuters. We then compare these results with those obtained under the three fare-setting scenarios examined earlier, and determine which scenario performs best under which conditions.
@@ 5.1 Formulation
The fare setting that achieves the optimal number of SAV commuters constitutes the second-best policy. The second-best fare, denoted by $p^{SB*}$, is obtained from the inverse demand function, derived by solving (27) for $p$:
@@
This price level is defined so that the mode-specific commuting costs satisfy $c_{n}^{*}=c_{a}^{*}$ when the number of SAV commuters is $N_{a}$. Substituting the second-best number of SAV commuters into this expression yields $p^{SB*}$. [^7] If the number of SAV commuters lies on a boundary (i.e., $N_{a}=0$ or $N_{a}=N$), the fare can be higher or lower than Eq. (64) because cost equality is no longer required, as shown in (13). To avoid unnecessary complications, we nevertheless use the price given by (64) in all cases.
@@ 5.2 SAV adoption and fare under the second-best policy
Setting $\partial Z^{SB}/\partial N_{a}=0$ yields the number of SAV commuters and the corresponding fare in the second-best equilibrium, stated in the following proposition.
@@
This proposition shows that the second-best SAV ridership and fare depend on the value of the index $\eta$. To clarify how they differ from the three fare-setting equilibria, we compare the second-best outcomes with the equilibrium SAV ridership and fare levels under each scenario. The comparison reveals that the ordering of these values changes depending on whether $\eta$ is greater than or equal to $1$. Specifically,
@@
    $\eta\geq 1$ $(\kappa\leq\theta)$: The second-best SAV ridership satisfies the following relationships with the equilibrium numbers of SAV commuters:
@@
This implies that, unlike the comparison of equilibrium commuting costs, MC pricing is not always the best policy for reducing the social cost. When $\eta\geq 1$, we have $1-\eta\leq 0$; Eq. (67) shows that the second-best fare is necessarily below the marginal cost $m$ and the corresponding SAV ridership exceeds the MC benchmark. This suggests that promoting SAV adoption through MC pricing reduces the social cost. By contrast, when $\eta<1$, we obtain $1-\eta>0$; hence, the second-best fare may exceed the marginal cost. In that case, MC pricing may induce excessive SAV adoption and the social cost can be higher than those under the other fare-setting scenarios.
@@
The first term is the marginal change in commuting cost incurred by NV commuters. The second and third terms describe the change in the total cost of SAV commuters. The second term captures how additional SAV commuters affect the congestion externalities they face, whereas the third term captures the change in the fare they pay, given that the mode-choice equilibrium is preserved. The fourth and fifth terms describe the change in the operator’s profit. The third and fifth terms offset each other because both represent income transfers effected through the fare.
@@
It is worth emphasizing that the additional externalities created when commuters shift from NVs to SAVs are converted into a reduction in the operator’s profit. In the second-best equilibrium, the fare is adjusted so that the mode-choice condition $c_{n}^{*}=c_{a}^{*}$ continues to hold. As a result, the increase in the congestion externalities for SAV commuters (second term) is offset by a corresponding decrease in the SAV fare (third term); this decrease causes the corresponding increase in the fifth term, which appears as a decrease in the operator’s revenue. Hence, although the increase in the social cost appears to stem from a decline in the operator’s revenue, attempting to restore that profit by further promoting SAV use would increase social cost even more. Therefore, reducing the social cost requires an appropriate price-regulation scheme that resolves the underlying problem of excessive SAV usage.
@@ 5.3 Social cost comparison under different fare-setting scenarios
Throughout this section, we assume that Eq. (23) holds strictly, so that SAVs are used under both AC pricing and the natural-monopoly fare. Specifically, we impose the following condition on the population size $N$:
@@
    $SC^{AC*}_{0}=\left[\cfrac{A}{1-\theta}N+t_{f}+F_{n}\right]N-F_{a},$ (72)
@@
    $SC^{m*}$ $=\left[\cfrac{(1+\kappa)AN+(1-\kappa)B}{2(1-\theta)}+t_{f}+F_{n}\right]N-\left[\cfrac{(AN-B)^{2}}{4A}+F_{a}\right].$ (74)
@@
where the critical population level $N^{MC=AC}_{c}$ is
$N^{MC=AC}_{c}:=\cfrac{\eta B+\sqrt{\eta^{2}B^{2}+4\eta(1-\eta)AF_{a}}}{2\eta(1-\eta)A}.$ (77)
@@
where the critical population level $N_{c}^{AC=m}$ is
$N_{c}^{AC=m}:=\cfrac{B+\sqrt{4\eta^{2}B^{2}+4(1-4\eta^{2})AF_{a}}}{(1-4\eta^{2})A}.$ (80)
@@
where the additional critical level $N_{c}^{MC=m}$ is given as follows:
$N_{c}^{MC=m}:=\cfrac{B}{(1-2\eta)A}.$ (82)
This proposition shows that when $\eta$ becomes very small, even the natural-monopoly scenario can outperform the other price regulations. Indeed, as $\eta$ approaches zero (i.e., $\kappa\rightarrow 1$), the following qualitative result holds.
@@
As $\eta\rightarrow 0$, the number of SAV commuters and the fare in the second-best equilibrium approach those under natural monopoly; that is,
@@
Figure 5: Contour plot of the social cost. $N=10,000$, $\theta=0.5$, $\beta=0.3$, $\gamma=2.0$, $\mu=1.0$, $t_{f}=2.0$, $F_{n}=10,000$, $F_{a}=1,000,000$, $m=10197$.
Figure 5 shows a contour plot of the social cost as a function of the number of SAV commuters, $N_{a}$, and the index $\eta$. In the plot, $\theta$ is held constant while $\kappa$ varies, so changes in $\eta$ arise solely from changes in the capacity-expansion parameter. First, when $\eta$ takes small values, the contours protrude toward the lower-left corner. This shape reveals that increasing $N_{a}$ does not necessarily reduce the social cost. For example, along the line $\eta=0.15$, the social cost initially reduces as $N_{a}$ increases, but begins to increase once the curve bends. As $\eta$ approaches zero, the socially optimal $N_{a}$ converges to the monopoly level $N_{a}^{m*}$. By contrast, when $\eta$ becomes large, the contours no longer protrude to the lower left. In this region, the capacity-expansion effect is strong, and increasing $N_{a}$ is socially desirable. Above a certain threshold for $\eta$, it is optimal for all commuters to use SAVs, as indicated by the closed-form solution for $N_{a}^{*}$.
5.4 Regulatory strategy lowering social cost while achieving a Pareto improvement
Based on the social cost analysis, we finally link the social cost to the regulatory strategy introduced in Section 3. We focus on the relationship between equilibrium commuting costs and social costs under AC pricing and natural monopoly. If $\eta\geq 1$ or $N<N_{c}^{AC=m}$, the two scenarios are ordered the same way for both objectives:

### v2

@@ 4 First-best policy
So far, we have examined how commuters’ equilibrium costs vary across three fare-setting scenarios. We now shift our attention to social costs that incorporate the service operator’s profit, and to transport policies designed to achieve an efficient transportation system.
This section examines the first-best policy that minimizes the social cost. Specifically, we assume that, in addition to fare regulation, a policymaker can impose time-varying congestion tolls that completely internalize congestion externalities. The equilibrium condition is formulated under the first-best policy. We then derive the equilibrium flow and cost patterns and analyze their properties.
@@ 4.1 Formulation
$\min_{\{n_{n}(t),n_{a}(t)\}_{t\in\mathbb{R}}}\quad$ $\begin{aligned} SC(\{n_{n}(t),n_{a}(t)\}):=\;&\int_{\mathbb{R}}\Big[n_{n}(t)\big(t_{f}+s(t)+F_{n}\big)\\ &\qquad\qquad+\,n_{a}(t)\big(\theta t_{f}+s(t)+m+w\big)\Big]\,\mathrm{d}t\;+\;F_{a}\end{aligned}$ (41)
s.t. $\int_{\mathbb{R}}n_{n}(t)\,\mathrm{d}t+\int_{\mathbb{R}}n_{a}(t)\,\mathrm{d}t=N,$ (42)
$n_{n}(t)+\kappa n_{a}(t)\leq\mu,\quad\forall t\in\mathbb{R},$ (43)
$n_{n}(t)\geq 0,\quad n_{a}(t)\geq 0,\quad\forall t\in\mathbb{R}.$ (44)
@@
We then consider a first-best policy that achieves this outcome in equilibrium. The policy consists of optimal fare regulation together with a time-varying congestion toll. Let $\tau(t)$ denote the congestion toll paid by NV commuters who arrive at their destination at time $t$. We assume that each SAV commuter pays $\kappa\tau(t)$, because an SAV imposes congestion externalities that are $\kappa$ times those generated by an NV commuter. [^9] Charging differentiated congestion tolls by vehicle class is technically feasible and has already been implemented in practice. An alternative is to establish, outside the transport system, a market for tradable access rights or credits that allows commuters to pre-purchase bottleneck permits priced according to their vehicle classes. We further assume that the SAV fare is set according to MC pricing: $p=m$. Under these assumptions, the commuting costs for NV and SAV commuters arriving at time $t$ are given by
@@
$c_{a}^{FB}(t)=s(t)+\theta t_{f}+m+w+\kappa\tau(t).$ (46)
The equilibrium conditions under the first-best policy are expressed as follows:
@@
Examining the optimality conditions of (41)–(44) shows that the proposed policy, i.e., MC pricing combined with vehicle-specific congestion tolls, achieves the first-best outcome. The equilibrium conditions coincide with the first-order conditions when the Lagrange multiplier for constraint (42) is set to $c^{FB*}$ and the multiplier for constraint (43) is set to $\tau(t)$. Consequently, an equilibrium that satisfies the equilibrium conditions attains the first-best optimum, and the proposed policy is therefore first best.
@@ 4.2 Characterization of the equilibrium under the first-best policy
The results indicate that a flow pattern consisting exclusively of SAVs cannot arise under the first-best equilibrium. The reason lies in the condition $B>0$, which represents that the marginal cost of using an SAV exceeds that of an NV. Because queues are eliminated in the first-best equilibrium, commuters’ mode choice depends only on differences in toll savings $(1-\kappa)\tau(t)$ and marginal costs $B$. During periods with low tolls (e.g., immediately after the bottleneck becomes active or just before it becomes inactive), the higher marginal cost of using an SAV outweighs the toll savings. As a result, commuters have an incentive to choose NVs.
An SAV becomes optimal when the toll savings exceed the marginal cost disadvantage, i.e., when $\tau(t)>B/(1-\kappa)$. Because the optimal toll is the shadow price of the capacity constraint, its peak level increases with the population size $N$: when $N$ is larger, the toll around the desired arrival time must increase to keep the bottleneck queue-free. Condition (46) gives the minimum $N$ such that the peak toll exceeds $B/(1-\kappa)$; when it holds, Case (i) arises and SAVs are used in the peak-neighborhood.
Figure 4: Time-varying congestion tolls in the first-best equilibrium. $N=10000$, $\theta=0.5$, $\beta=0.3$, $\gamma=2.0$, $\mu=1$, $t_{f}=2$, $F_{n}=500$, $m=800$, $w=200$.
Figure 4 illustrates the time-varying congestion tolls $\tau(t)$ for NV commuters and $\kappa\tau(t)$ for SAV commuters. In the figure, the tolls peak around the desired arrival time because SAVs are used during this period. An SAV commuter pays $\kappa\tau(t)$, which is $\kappa$ times the toll paid by an NV commuter and therefore lower. Taking this into account, the policymaker sets $\tau(t)$ sufficiently high to prevent queues from forming among SAV commuters. In addition, a comparison of $\tau(t)$ and $\kappa\tau(t)$ shows that their gap is large when SAVs are used and narrows as the destination arrival time moves away from the desired arrival time. Consequently, in periods where this gap is small, commuters have an incentive to choose NVs, whose marginal cost is lower than that of SAVs.
@@ 4.3 Pareto-improvement condition
Using the closed-form solution, we examine whether the first-best policy achieves a Pareto improvement, i.e., whether it reduces the social cost without increasing any commuter’s cost. To identify a sufficient condition, we compare the commuting cost $c^{FB*}$ in the first-best equilibrium with that under MC pricing $c^{MC*}$, which is the lowest commuting cost among the three toll-free fare-setting scenarios.
For this analysis, the following capacity-VOT index $\eta$, which summarizes the relative strength of capacity and VOT effects, plays a central role:
@@
Its value is determined by the relative magnitudes of $\kappa$ and $\theta$:
@@
This relationship shows that when the capacity effect dominates ($\eta>1$), $N_{a}^{FB*}$ exceeds $N_{a}^{MC*}$. It follows that the first-best policy requires an additional increase in the number of SAV commuters. As the SAV share increases, the total capacity effect grows. This reduces commuters’ schedule costs and hence their commuting costs.
By contrast, when $\eta<1$, the first-best policy reduces the number of SAV commuters. Consequently, the total capacity effect in the first-best equilibrium decreases, and schedule costs may increase. A similar phenomenon arises in Case (ii), where SAVs are not used: the first-best policy eliminates SAV usage, which increases commuting costs. These results suggest that congestion tolls should be introduced, taking into account the strength of the capacity effect, i.e., the technological maturity of SAVs.
@@ 4.4 Self-financing principle
We demonstrated that the first-best equilibrium can be achieved by combining MC pricing with a time-varying congestion toll. For a congestion-pricing scheme to be socially acceptable, it is desirable to redistribute the toll revenue to commuters. As a benchmark form of redistribution, we consider using the revenue to finance the expansion of bottleneck capacity. We then investigate whether the self-financing principle (Mohring and Harwitz, 1962) holds, namely, whether the revenue generated by the optimal congestion toll covers the cost of the optimal capacity level.
@@
$\begin{cases}\cfrac{\mathrm{d}K(\mu)}{\mathrm{d}\mu}-\int_{t\in\mathbb{R}}\tau(t)\mathrm{d}t=0\quad&\text{if}\quad\mu>0,\\ \cfrac{\mathrm{d}K(\mu)}{\mathrm{d}\mu}-\int_{t\in\mathbb{R}}\tau(t)\mathrm{d}t\geq 0\quad&\text{if}\quad\mu=0.\end{cases}$ (64)
@@ 5 Second-best policy
This section examines the second-best policy when congestion externalities cannot be internalized. We assume that the policymaker can regulate only the SAV fare and chooses it to minimize the social cost in equilibrium, subject to the departure-time and mode choice conditions. We first derive the second-best fare and the corresponding number of SAV commuters. We then compare these results with those under the three fare-setting scenarios examined earlier and identify which scenario yields the lowest social cost under which conditions.
@@ 5.1 Formulation
The second-best fare $p^{SB*}$ is chosen to achieve the optimal number of SAV commuters. It is obtained from the inverse demand function, which is derived by solving (27) for $p$:
@@
This inverse demand function gives the fare $p$ that satisfies the mode-choice indifference condition $c_{n}^{*}=c_{a}^{*}$ for a given $N_{a}$. Substituting $N_{a}=N_{a}^{SB*}$ into (69) yields $p^{SB*}$. [^10] If the number of SAV commuters lies on a boundary (i.e., $N_{a}=0$ or $N_{a}=N$), the fare can be higher or lower than Eq. (69) because cost equality is no longer required, as shown in (13). To avoid unnecessary complications, we nevertheless use the price given by (69) in all cases.
@@ 5.2 SAV adoption and fare under the second-best policy
Setting (70) to zero yields $N_{a}^{SB*}$ and the corresponding fare $p^{SB*}$, as stated in the following proposition:
@@
The proposition implies that $N_{a}^{SB*}$ and $p^{SB*}$ depend on the capacity-VOT index $\eta$. To see how the second-best outcomes differ from the equilibria under the three fare-setting scenarios, we compare $N_{a}^{SB*}$ with the corresponding equilibrium values of $N_{a}$ under each scenario. This reveals that the ordering of these values depends on whether $\eta\geq 1$ or not. Specifically,
@@
    $\eta\geq 1$ $(\kappa\leq\theta)$: The equilibrium values satisfies the following ordering:
@@
This implies that, unlike the comparison of equilibrium commuting costs, MC pricing is not always the best policy for reducing the social cost. When $\eta\geq 1$, we have $1-\eta\leq 0$. Eq. (72) shows that $p^{SB*}$ is below the marginal cost $m$, and $N_{a}^{SB*}$ exceeds $N_{a}^{MC*}$. In this case, MC pricing promotes SAV adoption toward the second-best level and yields a lower social cost than the other fare-setting scenarios. By contrast, when $\eta<1$, we obtain $1-\eta>0$; hence, $p^{SB*}$ may exceed $m$. This implies that MC pricing may induce excessive SAV adoption which leads to a higher social cost than that under the other fare-setting scenarios.
@@
The first term is the marginal change in commuting cost incurred by NV commuters. The second and third terms describe the change in the total cost of SAV commuters. The second term captures how additional SAV commuters affect the congestion externalities they face, whereas the third term captures the change in the fare they pay, given that the mode choice equilibrium is preserved. The fourth and fifth terms describe the change in the operator’s profit. The third and fifth terms offset each other because both represent income transfers effected through the fare.
@@
It is worth emphasizing that the additional externalities created when commuters shift from NVs to SAVs are converted into a reduction in the operator’s profit. In the second-best equilibrium, the fare is adjusted so that the mode choice condition $c_{n}^{*}=c_{a}^{*}$ continues to hold. As a result, the increase in the congestion externalities for SAV commuters (second term) is offset by a corresponding decrease in the SAV fare (third term); this decrease causes the corresponding increase in the fifth term, which appears as a decrease in the operator’s revenue. Hence, although the increase in the social cost appears to stem from a decline in the operator’s revenue, attempting to restore that profit by further promoting SAV use would increase social cost even more. Therefore, reducing the social cost requires an appropriate fare regulation scheme that resolves the underlying problem of excessive SAV usage.
@@ 5.3 Social cost comparison under different fare-setting scenarios
Throughout this section, we assume that Eq. (23) holds strictly, so that SAVs are used under both AC pricing and unregulated monopoly pricing. Specifically, we impose the following condition on the population size $N$:
@@
    $SC^{AC*}_{0}=\left[\cfrac{A}{1-\theta}N+t_{f}+F_{n}\right]N+F_{a},$ (77)
@@
    $SC^{m*}$ $=\left[\cfrac{(1+\kappa)AN+(1-\kappa)B}{2(1-\theta)}+t_{f}+F_{n}\right]N-\left[\cfrac{(AN-B)^{2}}{4A}-F_{a}\right].$ (79)
@@
where $N^{MC=AC}_{c}$ is the population threshold at which the ordering of social costs under MC and AC pricing reverses. [^13] The closed-form expressions of the population thresholds used throughout this section are also provided in Appendix F.
@@
where $N_{c}^{AC=m}$ is the population threshold at which the ordering of social costs under AC and unregulated monopoly pricing reverses.
@@
where $N_{c}^{MC=m}$ is the population threshold at which the ordering of social costs under MC and unregulated monopoly pricing reverses.
This proposition shows that, for sufficiently small $\eta$, the natural-monopoly scenario can yield the lowest social cost among the three fare-setting scenarios. Indeed, as $\eta\rightarrow 0$ (i.e., $\kappa\rightarrow 1$), the following qualitative result holds.
@@
As $\eta\rightarrow 0$, the number of SAV commuters and the fare in the second-best equilibrium approach those under unregulated monopoly pricing; that is,
@@
Figure 5: Contour plot of the social cost. $N=10000$, $\theta=0.5$, $\beta=0.3$, $\gamma=2.0$, $\mu=20$, $t_{f}=2$, $F_{n}=300$, $m=210$, $w=100$, $F_{a}=50000$.
Figure 5 shows a contour plot of the social cost as a function of the number of SAV commuters, $N_{a}$, and the index $\eta$. In the plot, $\theta$ is held constant while $\kappa$ varies, so changes in $\eta$ arise solely from changes in $\kappa$. First, when $\eta$ takes small values, the contours protrude toward the lower-left corner. This shape reveals that increasing $N_{a}$ does not necessarily reduce the social cost. For example, along the line $\eta=0.15$, the social cost initially decreases as $N_{a}$ increases, but begins to increase once the curve bends. As $\eta$ approaches zero, the socially optimal $N_{a}$ converges to the monopoly level $N_{a}^{m*}$. By contrast, when $\eta$ becomes large, the contours no longer protrude to the lower left. In this region, the capacity-expansion effect is strong, and increasing $N_{a}$ is socially desirable. Above a certain threshold for $\eta$, it is optimal for all commuters to use SAVs, as indicated by the closed-form solution for $N_{a}^{*}$.
5.4 Pricing strategy for social-cost reduction and Pareto improvement
Based on the social cost analysis, we finally link the social cost to the pricing strategy introduced in Section 3. We focus on the relationship between equilibrium commuting costs and social costs under AC pricing and unregulated monopoly pricing. If $\eta\geq 1$ or $N<N_{c}^{AC=m}$, AC pricing dominates unregulated monopoly pricing in terms of both commuters’ equilibrium cost and social cost: